# LatentExplainer: Explaining Latent Representations in Deep Generative Models with Multimodal Large Language Models

### v5

@@ 4.3 Automatic In-context Prompt Generation
4.4 Inductive-bias-guided Data Manipulation
@@
First, identify the relevant formulas within the input inductive bias formulas with regard to a specific latent variable $z_{i}$ to be explained. Then, combine the identified relevant formulas with the inductive bias prompt $\mathcal{P}$ obtained from Section 4.3 and utilize an LLM as a coding agent to generate prompts that specify the modifications needed in the generative model’s decoder code (e.g., adjusting indices through the subscripts of the latent variables or adding properties according to extra information about the symbol $\mathcal{I}$). The coding agent then tests and refines the generated code to effectively perturb $z_{i}$. Through a series of perturbations on $z_{i}$, a sequence of generated images can be obtained, capturing the variations in $z_{i}$ and reflecting its influence on the generated data. Finally, all image sequences generated from the relevant formulas are aggregated to generative explanations about $z_{i}$. The implementation details are in Appendix A.
4.5 Uncertainty-aware Explanations
@@
Uncertainty-aware methods can be applied to large language model responses Lin et al. (2023) and image explanations Zhao et al. (2024). To measure the uncertainty of the responses from GPT-4o, we sampled $n$ times from the GPT-4o to generate the responses =$\mathcal{R}=\{r_{1},r_{2},r_{3},...,r_{n}\}.$ The certainty score of the explanation is the average pairwise cosine similarity of the responses $\mathcal{R}$: $1/C\cdot{\sum}_{i=1}^{n}{\sum}_{j=1,i\neq j}^{n}\text{sim}(r_{i},r_{j})$ where $C=n\cdot(n-1)$. Our final explanation $\hat{r}$ is the response that has the highest pairwise cosine similarity with other responses if the latent variable is interpretable. Otherwise, $\hat{r}$ will be “No clear explanation”. The implementation details are in Appendix B.
5 Experiment
@@
The experiments are conducted on a 64-bit machine with 24-core Intel 13th Gen Core i9-13900K @ 5.80GHz, 32GB memory and NVIDIA GeForce RTX 4090. We use GPT-4o as our MLLM backbone set $temperature=1$ and $top\_p=1.$ The code is available at  https://anonymous.4open.science/r/LatentExplainer-89AF.
5.1 Dataset
@@
We utilized five datasets to evaluate the performance of different generative models under three inductive biases: CelebA-HQ Karras et al. (2018), AFHQ Choi et al. (2020), LSUN-Church Yu et al. (2015) for the unconditional and conditional diffusion models, and CelebA-HQ Karras et al. (2018), 3DShapes Burgess and Kim (2018), and dSprites Matthey et al. (2017) for the VAE models. The CelebA-HQ dataset is a high-quality version of the CelebA dataset, consisting of 30K images of celebrities, divided into 28K for training and 2K for testing. The LSUN-Church dataset contains large-scale images of church buildings. The AFHQ dataset includes high-quality images of animals, divided into three categories: cats, dogs, and wild animals. 3DShapes is a synthetic dataset contains images of 3D shapes with six factors of variation: floor hue, wall hue, object hue, object scale, object shape, and wall orientation, divided into 384K for training and 96K for testing. dSprites consists of 2D shapes (hearts, squares, ellipses) generated with five factors of variation: shape, scale, orientation, position X, and position Y, divided into 516K for training and 221K for testing.
@@
DDPM (Disentanglement Bias) Gemini 1.5 Pro 4.28 27.03 11.94 88.68 -3.33 5.38 24.35 8.88 87.45 -3.46 4.94 25.28 8.96 88.82 -3.08
Claude 3.5 Sonnet 12.98 28.7 18.68 89.24 -3.12 12.99 31.71 18.75 89.50 -3.25 13.17 31.62 14.99 89.51 -2.95
@@
Stable Diffusion (Conditional Bias) Gemini 1.5 Pro 0.00 20.52 12.99 88.87 -3.31 2.29 21.53 8.15 88.08 -3.53 0.00 21.46 7.51 88.43 -3.11
Claude 3.5 Sonnet 6.43 29.90 17.54 89.59 -3.24 0.00 28.53 13.8 88.85 -3.47 4.22 25.71 13.63 89.20 -3.10
@@
Table 2: Quantitative results for diffusion models across datasets. B represents BLEU@4, R represents ROUGE-L, S represents SPICE, BS represents BERTScore, and BAS represents BARTScore.
@@
$\beta$-TCVAE (Disentanglement Bias) Gemini 1.5 Pro 9.31 31.31 8.63 89.50 -3.12 4.56 35.10 12.82 89.87 -3.11 0.98 16.47 7.67 86.39 -3.25
Claude 3.5 Sonnet 10.37 31.39 13.3 89.13 -2.94 11.81 32.89 13.08 89.92 -3.01 5.45 29.72 15.54 89.91 -2.97
@@
CSVAE (Combination Bias) Gemini 1.5 Pro 6.86 25.14 5.88 89.08 -2.98 8.41 24.84 12.49 89.36 -3.09 0.00 16.86 5.6 87.17 -3.21
Claude 3.5 Sonnet 16.11 28.98 21.05 89.05 -2.81 17.98 28.61 20.73 89.50 -2.95 14.28 27.58 18.98 $8$2.94
@@
CSVAE (Conditional Bias) Gemini 1.5 Pro 9.94 28.88 8.06 89.36 -3.10 2.42 17.96 7.02 88.02 -3.10 3.64 25.51 10.89 89.27 -3.10
Claude 3.5 Sonnet 9.00 33.02 15.69 88.48 -2.93 11.75 32.45 19.10 90.68 -2.93 0.00 26.21 16.88 88.62 -3.07
@@
Table 3: Quantitative results for VAE models across datasets. B represents BLEU@4, R represents ROUGE-L, S represents SPICE, BS represents BERTScore, and BAS represents BARTScore.
5.2 Models and Baselines
@@
Our evaluation benchmarks our proposed LatentExplainer framework against three state-of-the-art multimodal models with strong vision-language reasoning capabilities: GPT-4o OpenAI (2024), Gemini 1.5 Pro Team et al. (2024), and Claude 3.5 Sonnet Anthropic (2024). We employ GPT-4o as a zero-shot baseline, comparing it with the addition of the inductive bias prompt, uncertainty quantification, and the full model with both components included.
5.3 Generative Models under Inductive Biases
@@
We explore the latent space in generative models that satisfy the aforementioned three types of inductive biases. For each type, we present the relevant generative models that align with the corresponding inductive bias: (1) Disentanglement Bias: $\beta$-TCVAE Chen et al. (2018) explicitly penalizes the total correlation of the latent variables to disentangle the latent representations. Denoising Diffusion Probabilistic Model (DDPM) Ho et al. (2020) adds Gaussian noise independently at each timestep in the forward process and eventually transforms into pure Gaussian noise, in which the covariance matrix is diagonal. This assumes the latent factors are independent; (2) Combination Bias: CSVAE Klys et al. (2018) has two groups of latent variables $z$ and $w$, where $z$ and $w$ are uncorrelated and the latent variables within the group are correlated; (3) Conditional Bias: CSVAE also satisfies conditional bias because one group of latent variables $w$ is associated with the properties while the other group of latent variables $z$ minimizes the mutual information with the properties. Stable Diffusion Rombach et al. (2022) is a latent diffusion model to generate images conditioned on prompts.
@@
Figure 3: Visualization of the generated explanations with the inductive bias prompt for the disentanglement bias.
@@
Figure 4: Visualization of the generated explanations w/o the inductive bias prompt for the disentanglement bias.
5.4 Quantitative Analysis
For the quantitative explanation evaluation, we use BLEU Papineni et al. (2002), ROUGE-L Lin (2004), SPICE Anderson et al. (2016), BERTScoreZhang et al. (2019), and BARTScore Yuan et al. (2021) as the automated metrics to assess the generated explanations. BLEU, and ROUGE-L are n-gram-based metrics that measure the overlap between generated and reference texts. SPICE compares scene graphs derived from the generated and reference texts. BERTScore and BARTScore utilize pre-trained transformer-based language models to compute contextual embeddings of the generated and reference texts. These metrics together provide a comprehensive assessment of both the lexical and semantic quality of the explanations.
To understand the impact of various components in our proposed LatentExplainer, we conducted a comprehensive quantitative analysis across different models (DDPM, $\beta$-TCVAE, Stable Diffusion, and CSVAE) and inductive biases (disentanglement, combination, and conditional). Specifically, we compare the removal of inductive bias prompts (IB), the removal of uncertainty quantification (UQ), and the full model against the baseline GPT-4o. The results for each dataset are provided in Table 2 and Table 3. Removing inductive bias prompts leads to a substantial drop in all generative models. Their consistent results demonstrate that inductive bias is the most important and necessary component when explaining the latent representations of generative models. The removal of uncertainty quantification also results in a slightly decreased performance in all generative models, indicating that uncertainty quantification is also effective, though not as critical as inductive bias prompts. The full model, which incorporates both inductive bias prompts and uncertainty quantification, achieves the highest overall performance, and outperforms all baselines across all models. This confirms the necessity of both inductive bias prompts and uncertainty quantification in our LatentExplainer framework, demonstrating their significant contributions to improving explanatory performance across various generative models. It also shows that our framework can effectively design prompts for different inductive biases in generative models to improve the accuracy and reduce hallucination.
In addition to GPT-4o, we evaluate the explanation performance of Gemini 1.5 Pro and Claude 3.5 Sonnet across all generative models and datasets. Both Gemini 1.5 Pro and Claude 3.5 Sonnet underperform compared to our LatentExplainer approach. Our method demonstrates substantial gains in all diffusion and VAE models, particularly for BLEU@4, ROUGE-L, and SPICE, suggesting that explicitly verbalizing inductive biases significantly enhances explanation quality. Furthermore, while Claude 3.5 Sonnet achieves high semantic quality, it does not exhibit the same level of consistency across datasets and models as our proposed method. Overall, while GPT-4o, Claude 3.5 Sonnet and Gemini 1.5 Pro provide competitive baselines, our LatentExplainer consistently improves both lexical and semantic quality, demonstrating its effectiveness in enhancing explainability for latent variables.
5.5 Qualitative Evaluation
To analyze the explanations for DDPM under disentanglement bias, we manipulate the latent representation along a latent direction and compare it with the one that first traverses along another latent direction and then traverses along the same latent direction. The disentangled latent variable would be invariant with respect to the variations in another latent dimension. In Figure 4, for each latent direction, we pass two image sequences and an inductive bias prompt based on the disentanglement formula to the MLLM to obtain a common latent explanation. In comparison, the explanations generated without the inductive bias prompt as shown in Figure 4 tend to show “no clear explanation” or wrong explanations as they only align with one image sequence but do not reflect the common pattern in both image sequences in view of the inductive bias. The addition of the inductive bias prompts can assist with ruling out the variation effects of other latent variables to capture the actual meaning.
We also qualitatively evaluate the explanations for CSVAE under combination bias. We transverse a latent variable and compare it with the one that first traverses another latent variable in another group and then traverses the same latent variable, which is similar to the disentanglement bias. We then compare with the one that first transverses another latent variable in the same group and then traverses the same latent variable. As Figure 10 depicts, our model can clearly show the explanations of latent variable $z_{1},z_{2},z_{3}$ as the color of the ground, the background color, and the shape of the object. The effect of removing inductive bias prompts leads to no clear explanation or wrong explanations like the disentanglement bias.
In Figure 11, we provide the “young appearance” prompt to Stable Diffusion under conditional bias, and the explanations of all three top latent directions reflect the meaning of youth. The addition of inductive bias prompts can better identify the relation with the property of interest to capture the actual meaning of latent variables. In comparison, the one without the inductive bias prompts in Figure 12, cannot find clear explanations or simply describe the characteristics in the image sequence, lacking an abstract generalization. More qualitative evaluation results can be found in Appendix E.
6 Conclusion
Appendix A Latent Variable Perturbation
@@
For VAE models, every time we target at each latent variable. The targeted latent variable is changed from -3 to 3, which is the 3 times the standard deviation of the standard Gaussian distribution of a latent variable, with a step size of 1.5. All the other latent variables are unchanged. The value of a specific latent variable among other latent variables can be determined by decomposing the inductive bias formulas. For diffusion models, we follow $\tilde{\mathbf{z}}=\mathbf{z}+\gamma\left[\mathcal{G}(\mathbf{z}+\mathbf{v})-\mathcal{G}(\mathbf{z})\right]$, where $\gamma$ is the strength and $\mathcal{G}$ is a diffusion decoder. We set the value of $\gamma$ to 0.1, 0.2, 0.3, 0.4, and 0.5 for DDPM, and to 1, 2, 3, 4, and 5 for Stable Diffusion. $\mathbf{v}$ may be determined by decomposing the inductive bias formulas.
Appendix B Uncertainty-aware Implementations

### v6

@@ 4.3. Automatic In-context Prompt Generation
4.4. Inductive-bias-guided Data Manipulation
@@
First, identify the relevant formulas within the input inductive bias formulas with regard to a specific latent variable $z_{i}$ to be explained. Then, combine the identified relevant formulas with the inductive bias prompt $\mathcal{P}$ obtained from Section 4.3 and utilize an LLM as a coding agent to generate prompts that specify the modifications needed in the generative model’s decoder code (e.g., adjusting indices through the subscripts of the latent variables or adding properties according to extra information about the symbol $\mathcal{I}$). The coding agent then tests and refines the generated code to effectively perturb $z_{i}$. Through a series of perturbations on $z_{i}$, a sequence of generated images can be obtained, capturing the variations in $z_{i}$ and reflecting its influence on the generated data. Finally, all image sequences generated from the relevant formulas are aggregated to generate explanations about $z_{i}$. The implementation details are in Section 5.4.
4.5. Uncertainty-aware Explanations
@@
Uncertainty-aware methods can be applied to large language model responses (Lin et al., 2023) and image explanations (Zhao et al., 2024). To measure the uncertainty of the responses from GPT-4o, we sampled $n$ times from the GPT-4o to generate the responses =$\mathcal{R}=\{r_{1},r_{2},r_{3},...,r_{n}\}.$ The certainty score of the explanation is the average pairwise cosine similarity of the responses $\mathcal{R}$: $1/C\cdot{\sum}_{i=1}^{n}{\sum}_{j=1,i\neq j}^{n}\text{sim}(r_{i},r_{j})$ where $C=n\cdot(n-1)$. Our final explanation $\hat{r}$ is the response that has the highest pairwise cosine similarity with other responses if the latent variable is interpretable. Otherwise, $\hat{r}$ will be “No clear explanation”. The implementation details are in Section 5.5.
5. Experiment
@@
The experiments are conducted on a 64-bit machine with 24-core Intel 13th Gen Core i9-13900K @ 5.80GHz, 32GB memory and NVIDIA GeForce RTX 4090. We set GPT-4o at $temperature=1$ and $top\_p=1.$ The code is available at  https://github.com/mengdanzhu/LatentExplainer.
5.1. Dataset
@@
We utilized five datasets to evaluate the performance of different generative models under three inductive biases: CelebA-HQ (Karras et al., 2018), AFHQ (Choi et al., 2020), LSUN-Church (Yu et al., 2015) for the unconditional and conditional diffusion models, and CelebA-HQ (Karras et al., 2018), 3DShapes (Burgess and Kim, 2018), and dSprites (Matthey et al., 2017) for the VAE models. The CelebA-HQ dataset is a high-quality version of the CelebA dataset, consisting of 30K images of celebrities, divided into 28K for training and 2K for testing. The LSUN-Church dataset contains large-scale images of church buildings. The AFHQ dataset includes high-quality images of animals, divided into three categories: cats, dogs, and wild animals. 3DShapes is a synthetic dataset contains images of 3D shapes with six factors of variation: floor hue, wall hue, object hue, object scale, object shape, and wall orientation, divided into 384K for training and 96K for testing. dSprites consists of 2D shapes (hearts, squares, ellipses) generated with five factors of variation: shape, scale, orientation, position X, and position Y, divided into 516K for training and 221K for testing.
@@
DDPM (Disentanglement Bias) Gemini 1.5 Pro 2.87 23.18 11.46 88.27 -3.35 1.64 20.75 6.92 87.22 -3.51 1.80 22.18 8.45 87.96 -3.11
+ LatentExplainer 3.15 25.62 13.36 88.63 -3.30 5.71 26.71 10.31 87.95 -3.46 4.97 26.74 9.64 88.72 -3.08
Claude 3.5 Sonnet 4.16 22.25 13.34 88.02 -3.28 7.76 28.02 14.14 88.65 -3.36 8.05 25.60 12.11 88.46 -3.10
+ LatentExplainer 12.52 31.68 20.05 89.75 -3.06 12.92 32.55 19.52 89.70 -3.23 14.14 34.07 14.80 90.06 -2.94
@@
Stable Diffusion (Conditional Bias) Gemini 1.5 Pro 0.00 20.36 10.95 88.74 -3.34 0.00 20.24 8.98 88.08 -3.56 0.00 18.93 7.48 87.83 -3.16
+ LatentExplainer 3.29 21.52 15.21 89.09 -3.29 0.00 21.49 9.57 88.09 -3.53 1.89 21.22 8.60 88.01 -3.09
Claude 3.5 Sonnet 6.57 30.62 17.80 89.53 -3.31 2.37 26.95 13.10 88.74 -3.55 4.69 26.76 13.52 89.27 -3.15
+ LatentExplainer 7.69 31.84 18.27 89.85 -3.28 2.69 28.88 14.14 89.01 -3.46 4.02 28.68 15.22 89.61 -3.08
Table 2. Quantitative results for diffusion models across datasets. B represents BLEU@4, R represents ROUGE-L, S represents SPICE, BS represents BERTScore, and BAS represents BARTScore.
Model Method 3DShapes CelebA-HQ dSprites
[…]
$\beta$-TCVAE (Disentanglement Bias) Gemini 1.5 Pro 2.73 26.45 5.56 88.83 -3.15 0.00 17.28 4.86 87.00 -3.27 0.00 15.39 6.11 86.61 -3.25
+ LatentExplainer 3.11 28.98 8.07 88.94 -3.14 2.19 31.21 10.70 88.98 -3.18 1.54 17.29 7.61 86.98 -3.22
Claude 3.5 Sonnet 11.02 30.95 14.13 89.11 -3.01 9.54 30.96 11.11 88.76 -3.14 3.65 22.32 11.91 88.49 -3.05
+ LatentExplainer 18.09 33.69 14.26 89.36 -2.92 12.73 33.27 11.21 89.50 -3.05 12.20 35.87 19.29 90.58 -2.95
CSVAE (Combination Bias) Gemini 1.5 Pro 3.11 22.14 5.39 88.33 -2.99 1.21 16.52 5.28 88.25 -3.21 1.18 23.15 5.83 88.24 -3.15
+ LatentExplainer 6.22 25.59 5.69 88.98 -2.97 3.88 24.17 7.94 88.72 -3.21 1.33 26.43 5.88 89.18 -3.08
Claude 3.5 Sonnet 4.55 24.30 9.14 87.08 -3.01 9.43 28.77 18.31 89.50 -3.11 7.04 24.16 14.27 88.45 -3.08
+ LatentExplainer 15.24 25.57 19.82 88.72 -2.87 19.92 28.85 18.97 89.55 -2.97 17.75 30.87 22.14 89.58 -2.94
CSVAE (Conditional Bias) Gemini 1.5 Pro 8.61 26.41 7.11 87.95 -3.20 3.48 19.77 10.41 87.92 -3.12 0.00 17.27 8.65 86.80 -3.30
+ LatentExplainer 12.12 35.80 9.38 90.14 -3.08 3.63 21.26 10.65 87.95 -3.05 2.62 22.99 10.53 88.26 -3.18
Claude 3.5 Sonnet 5.36 26.69 15.28 88.75 -3.04 13.83 34.38 18.15 90.36 -2.98 8.16 25.09 17.70 88.18 -3.11
+ LatentExplainer 8.42 31.32 18.34 88.82 -2.93 14.53 34.54 19.58 90.70 -2.90 8.42 27.99 17.94 88.34 -3.04
Table 3. Quantitative results for VAE models across datasets. B represents BLEU@4, R represents ROUGE-L, S represents SPICE, BS represents BERTScore, and BAS represents BARTScore.
Model Method AFHQ LSUN-Church CelebA-HQ
[…]
DDPM (Disentanglement Bias) GPT-4o 3.08 14.83 8.78 87.55 -3.39 5.52 23.06 11.23 88.47 -3.41 0.11 3.54 1.22 85.97 -3.16
@@
Stable Diffusion (Conditional Bias) GPT-4o 7.61 26.32 17.19 90.04 -3.32 7.80 26.73 14.28 89.44 -3.46 5.79 23.89 12.68 89.63 -3.04
@@
Table 4. Ablation results for diffusion models across datasets on GPT-4o. B represents BLEU@4, R represents ROUGE-L, S represents SPICE, BS represents BERTScore, and BAS represents BARTScore.
@@
$\beta$-TCVAE (Disentanglement Bias) GPT-4o 5.51 29.77 7.92 89.18 -3.13 11.09 37.14 17.24 90.55 -3.07 0.00 16.19 7.04 87.17 -3.22
@@
CSVAE (Combination Bias) GPT-4o 6.93 25.07 9.70 89.42 -2.87 10.44 37.42 15.06 90.53 -3.01 0.00 21.61 6.99 87.19 -3.17
@@
CSVAE (Conditional Bias) GPT-4o 10.97 30.55 8.83 90.10 -3.02 8.46 28.78 15.78 89.62 -2.98 0.00 9.62 1.03 85.74 -3.34
@@
Table 5. Ablation results for VAE models across datasets on GPT-4o. B represents BLEU@4, R represents ROUGE-L, S represents SPICE, BS represents BERTScore, and BAS represents BARTScore.
5.2. Models and Baselines
@@
Our evaluation benchmarks our proposed LatentExplainer framework against three state-of-the-art multimodal models with strong vision-language reasoning capabilities: GPT-4o (OpenAI, 2024), Gemini 1.5 Pro (Team et al., 2024), and Claude 3.5 Sonnet (Anthropic, 2024). We employ GPT-4o, Gemini 1.5 Pro, and Claude 3.5 as a zero-shot baseline, comparing it with the addition of our Latent Exlainer with both the inductive bias prompt and uncertainty quantification included.
5.3. Generative Models under Inductive Biases
@@
We explore the latent space in generative models that satisfy the aforementioned three types of inductive biases. For each type, we present the relevant generative models that align with the corresponding inductive bias: (1) Disentanglement Bias: $\beta$-TCVAE (Chen et al., 2018) explicitly penalizes the total correlation of the latent variables to disentangle the latent variables. Denoising Diffusion Probabilistic Model (DDPM) (Ho et al., 2020) adds Gaussian noise independently at each timestep in the forward process and eventually transforms into pure Gaussian noise, in which the covariance matrix is diagonal. This assumes the latent factors are independent; (2) Combination Bias: CSVAE (Klys et al., 2018) has two groups of latent variables $z$ and $w$, where $z$ and $w$ are uncorrelated and the latent variables within the group are correlated; (3) Conditional Bias: CSVAE also satisfies conditional bias because one group of latent variables $w$ is associated with the properties while the other group of latent variables $z$ minimizes the mutual information with the properties. Stable Diffusion (Rombach et al., 2022) is a latent diffusion model to generate images conditioned on prompts.
@@
Figure 3. Visualization of the generated explanations with the inductive bias prompt for the disentanglement bias.
@@
Figure 4. Visualization of the generated explanations w/o the inductive bias prompt for the disentanglement bias.
Figure 5. Visualization of the generated explanations with the inductive bias prompt for the combination bias.
Figure 6. Visualization of the generated explanations without the inductive bias prompt for the combination bias.
Figure 7. Visualization of the generated explanations with the inductive bias prompt for the conditional bias.
Figure 8. Visualization of the generated explanations without the inductive bias prompt for the conditional bias.
5.4. Latent Variable Perturbation
@@
For VAE models, every time we target at each latent variable. The targeted latent variable is changed from -3 to 3, which is the 3 times the standard deviation of the standard Gaussian distribution of a latent variable, with a step size of 1.5. All the other latent variables are unchanged. The value of a specific latent variable among other latent variables can be determined by decomposing the inductive bias formulas. For diffusion models, we follow $\tilde{\mathbf{z}}=\mathbf{z}+\gamma\left[\mathcal{G}(\mathbf{z}+\mathbf{z_{i}})-\mathcal{G}(\mathbf{z})\right]$, where $\gamma$ is the strength and $\mathcal{G}$ is a diffusion decoder. We set the value of $\gamma$ to 0.1, 0.2, 0.3, 0.4, and 0.5 for DDPM, and to 1, 2, 3, 4, and 5 for Stable Diffusion. $\mathbf{z_{i}}$ is determined by decomposing the inductive bias formulas.
5.5. Uncertainty-aware Implementations